if it takes 2 people 3 days to do a job how long would it take 5 people to do it
e.mail bruce@proshoots.com

Found 2 solutions by flame8855, slaborn:
Answer by flame8855(424)

w/2= 3 w=6
w/5=6/5=1.2 days

Answer by slaborn(1)

Days per man to complete a job = 3 days * 2 people = 6
Days = 6 / Number of people
Days = 6 / 5 = 1.2 days